R. Bantsuri

A. Razmadze Mathematical Institute, Georgian Academy of Sciences Tbilisi, Georgia

APPLICATION OF THE METHODS OF THE THEORY OF ANALYTIC FUNCTIONS IN THE THEORY OF ELASTICITY

The paper gives a survey of the results obtained for the last decade by the Georgian scientists on the problems of elasticity with the use of the theory of analytic functions. A description of basic stages of the development of this area is presented.